### v2

# NEDS-SLAM: A Novel Neural Explicit Dense Semantic SLAM Framework using 3D Gaussian Splatting

Abstract
We propose NEDS-SLAM, an Explicit Dense semantic SLAM system based on 3D Gaussian representation, that enables robust 3D semantic mapping, accurate camera tracking, and high-quality rendering in real-time. In the system, we propose a Spatially Consistent Feature Fusion model to reduce the effect of erroneous estimates from pre-trained segmentation head on semantic reconstruction, achieving robust 3D semantic Gaussian mapping. Additionally, we employ a lightweight encoder-decoder to compress the high-dimensional semantic features into a compact 3D Gaussian representation, mitigating the burden of excessive memory consumption. Furthermore, we leverage the advantage of 3D Gaussian splatting, which enables efficient and differentiable novel view rendering, and propose a Virtual Camera View Pruning method to eliminate outlier GS points, thereby effectively enhancing the quality of scene representations. Our NEDS-SLAM method demonstrates competitive performance over existing dense semantic SLAM methods in terms of mapping and tracking accuracy on Replica and ScanNet datasets, while also showing excellent capabilities in 3D dense semantic mapping.
Index Terms:
3D Gaussian Splatting; Dense Semantic Mapping; Neural SLAM; 3D Reconstruction.
I Introduction
Visual SLAM (Simultaneous Localization and Mapping) is a fundamental research problem in robotics, which involves simultaneously tracking the camera pose and incrementally constructing a map of an unknown environment [1]. Downstream tasks such as autonomous goal navigation, human-computer interaction, mixed reality (MR), and augmented reality (AR) demand not only accurate camera pose tracking from SLAM systems but also robust and dense semantic reconstruction of the environment. This research focuses on semantic RGBD-SLAM, which, in contrast to traditional SLAM, enables the identification, classification, and association of entities within a scene, ultimately generating a semantically-rich map.
Inspired by the success of NERF and 3D Gaussian Splatting (3DGS) in high-fidelity view synthesis, researchers have explored building end-to-end visual SLAM systems based on neural radiance fields. These novel SLAM architectures offer superior solutions compared to traditional algorithms in terms of surface continuity, memory requirements, and scene completion. Specifically, iMAP [2] and NICE-SLAM [3] leverage neural implicit fields for consistent geometry representation, while MonoSLAM and SplatAM employ 3D Gaussian Splatting (3DGS) to achieve photo-realistic mapping.
Existing radiance field-based SLAM methods primarily focus on RGB reconstruction, while only a few approaches, such as SGS-SLAM [4] and DNS-SLAM [5], address semantic reconstruction. However, these methods rely on accurate and consistent semantic pre-segmentation, and their reconstruction performance suffers significantly from any substantial deviations in semantic feature estimation. In practical SLAM scenarios, the inconsistency of semantic feature estimation is a non-negligible issue.
Overall, Neural Explicit Dense Semantic SLAM can be summarized as facing two key challenges: 1) Providing robust semantic reconstruction results under inconsistent semantic features. 2) Incrementally building a map that can accurately distinguish well-optimized and low-quality regions, while effectively filtering out outliers to improve reconstruction quality.
This paper proposes NEDS-SLAM, with the following key contributions:
  • We propose a fusion module that combines semantic features with appearance features, addressing the spatially inconsistency of semantic features and providing a more robust semantic SLAM solution.
  • We build a semantic SLAM framework based on 3DGS, embedding semantic features with a lightweight encoder-decoder to achieve accurate semantic reconstruction and photo-realistic reconstruction.
  • We introduce a virtual camera view pruning method to remove noisy Gaussians, enabling more accurate construction of the 3D Gaussian radiance field.
II Related Work
II-A Traditional approaches to dense semantic SLAM
Semantic information is paramount for SLAM systems, as it enables scene understanding beyond geometric reconstruction alone. The ability to perceive and model semantics represents a crucial requirement for enabling applications of SLAM in robotics, virtual or augmented reality, and other domains that require understanding of the environment [6],[7]. Real-time dense semantic SLAM systems face the challenge of effectively fusing semantic information into underlying 3D geometric representations of the environment. Traditional approaches using voxels, point clouds, and signed distance fields to encode object labels [8],[9]. However, voxel- and point cloud-based approaches struggle with reconstruction speed and high-fidelity model acquisition. Meanwhile, signed distance field representations incur high memory usage that does not scale well to large-scale environments. There remains a need for more efficient and expressive 3D semantic modeling techniques suitable for real-time dense SLAM.
II-B Gaussian Splatting based SLAM
3D Gaussian representations have emerged as a promising approach for 3D scene modelling using a set of 3D Gaussians, each characterized by parameters such as position, anisotropic covariance, opacity, and color [10]. While existing GS-based SLAM methods have primarily focused on RGB reconstruction, exploring end-to-end system architectures, optimization of GS parameters, and accurate camera pose tracking through differentiable rendering, less attention has been paid to semantic reconstruction [11],[12],[13],[14] The few semantic GS SLAM approaches proposed to date have simply encoded ground truth semantic color labels directly as a second color channel of the GS parameters [4], without explicit modeling of semantic information or inference. There is clear potential for more sophisticated integration of semantics within the GS framework. The present work conducts a more in-depth exploration of semantic GS SLAM, aiming to simultaneously improve the robustness and reconstruction fidelity of GS-based SLAM systems through more sophisticated modeling and inference of semantic information within the GS representation.
Figure 1: Overview of the proposed NEDS-SLAM. Our method takes an RGB-D stream as input. RGB images are processed by the pretrained semantic feature extractor to get semantic features, while dense appearance features are obtained through the Spatial Feature Extractor model. The semantic and appearance features are fused to generate high-dimensional semantic features that are spatially consistent. These features are then processed by the encoder to generate low-dimensional features and embedded into the GS parameters. By employing Differentiable Rendering, real RGB images, depth images, and semantic masks predicted by a pre-trained segmentation head are utilized for Multi-Channel supervision. This approach enables the joint optimization of GS parameters. In the figure, $M$, $C$, and $D$ represent the semantic segmentation mask, color, and depth information, respectively. NEDS-SLAM achieves high-fidelity map reconstructions while simultaneously accomplishing compact and dense pixel-level semantic reconstruction.
III Methodology
III-A Scene Representation and Semantic embedding
Each 3DGS utilized for representing three-dimensional scenes encompasses mean, covariance, and color information. In this paper, a simplified 3DGS representation of the scene is employed [11], omitting the spherical harmonics functions used for color representation, while assuming GS to be isotropic.
$f^{gs}\left(\mathbf{x}\right)=o\exp\left({-\frac{{{{\left\|{\mathbf{x}-\mu}\right\|}^{2}}}}{{2{r^{2}}}}}\right)$ (1)
Where $\mu\in\mathbb{R}^{3}$ represents the center position of the GS, $r$ is the radius, and $o\in[0,1]$ represents the opacity. The rapid and differentiable rendering based on GS splatting serves as the core of mapping and tracking within GS-based SLAM systems. This ability for fast rendering enables the system to directly compute the gradients of the underlying GS parameters based on the discrepancy between the rendered results and the actual data. Consequently, the GS parameters can be updated to achieve an accurate representation of the scene. The differentiable rendering process based on GS splatting comprises three steps: Frustum Culling, Splatting, and Rendering by Pixels [15].
$C\left(p\right)=\sum\limits_{i\in N}{\mathbf{c}_{i}}f_{i}^{gs}\left(p\right)\prod\limits_{j=1}^{i-1}{\left({1-f_{j}^{gs}\left(p\right)}\right)}$ (2)
After arranging a collection of 3D Gaussians and camera pose, it is imperative to sort the Gaussians in a front-to-back manner. By employing alpha-compositing, the splatted 2D projection of each Gaussian can be efficiently rendered in pixel space, ensuring the generation of RGB images in the desired order, as Eq. (2). $\mathbf{c}_{i}$ represents the color parameters of the GS, and $f_{i}^{gs}\left(p\right)$ is computed as in Eq. (1) but with the 2D splatted ${\mu}$ and ${r}$. The rendering process is completed by multiplying the opacity of each GS with the color and accumulating the results. The depth map is rendered in a similar manner, as shown in Eq. (3).
$D\left(p\right)=\sum\limits_{i\in N}{\mathbf{d}_{i}}f_{i}^{gs}\left(p\right)\prod\limits_{j=1}^{i-1}{\left({1-f_{j}^{gs}\left(p\right)}\right)}$ (3)
The most notable distinction between semantic features and color and geometric features lies in their high-dimensional attributes. The semantic features do not refer to the per-pixel class labels generated by the segmentation head. Instead, it pertains to the high-dimensional semantic features extracted by the pre-trained model at each pixel. Taking DINO [16] as an example, the ViT-S model produces latent feature encodings of 384 dimensions, while the ViT-G model produces encodings of 1536 dimensions.
A simple way to combine 3D Gaussian Splatting with semantic features is to add trainable feature vectors to each Gaussian distribution. These parameters can be learned during the differentiable rendering process, which allows end-to-end training. However, for explicit Dense Semantic SLAM, adding a high dimensional semantic feature vector to each 3DGS is memory-inefficient. It would also weaken the real-time performance of optimizing model parameters. Inspired by LangSplat [17], we propose using a simple MLP as an encoder to compact high-dimensional semantic features into a low-dimensional vector. The compressed semantic features are then added to the 3D gaussian splats and can be rendered as in Eq. (4).
$S\left(p\right)=\sum\limits_{i\in N}{\mathbf{f}_{i}}f_{i}^{gs}\left(p\right)\prod\limits_{j=1}^{i-1}{\left({1-f_{j}^{gs}\left(p\right)}\right)}$ (4)
III-B Adaptive 3D Gaussian Expansion Mapping
III-B1 Spatially Consistent Feature Fusion
Consistent and continuous semantic labels are crucial for 3D semantic mapping. Semantic SLAM uses pretrained semantic segmentation models to compute pixel-level semantic labels from each RGB frame, but these class labels lack environmental specificity. Pretrained models may produce inconsistent semantic estimates, where the same object is predicted with different semantic labels in images from different angles. This would significantly reduce the quality of constructing 3D semantic Gaussians.
Figure 2: Using Mask2Former [18] for semantic segmentation on the Replica dataset, there is a noticeable inconsistency in semantic labels between frame 0 and frame 30 (highlighted in white boxes in the image). The floor and chairs exhibit significant differences, and directly utilizing the semantic distribution of the scene for GS construction would result in incompact semantics.
To address this issue, SNI-SLAM [19] computes a fused feature by combining geometry, appearance, and semantic features, replacing the reliance on a single semantic label. CoSSegGaussians [20] incorporates DINO [16] features with superior multi-view semantic scale consistency into the GS parameters. Subsequently, the semantic encoding of each GS is fused with spatial coordinates to render semantic features, thereby enhancing robustness. Inspired by SNI-SLAM and CoSSegGaussians, this paper proposes a simplified fusion mechanism. It combines the appearance features estimated by DepthAnything [21] with the semantic features extracted from pretrained model. The resulting mixed feature, obtained through MLP computations, is then embedded as the final semantic encoding in the 3DGS representations.
Due to limitations imposed by sensor performance and real-world noise, the depth map is not densely populated at the pixel level. The utilization of DepthAnything as spatial feature extractor to estimate the relative depth map of the current frame aims to address the issue of holes and artifacts in the depth observation. Within the SLAM pipeline, an effective mask is computed for each depth frame, rendering pixels without depth information as masked out. These masked pixels are unable to contribute to the establishment and optimization of subsequent GS, even if they possess accurate RGB and semantic values. By employing DepthAnything, it ensures that the geometric information of the entire scene within the field of view is incorporated into the extraction of semantic features and the modeling process of semantic GS.
The relative depth between pixels can reflect the geometric structure of observed object surfaces. The feature fusion model dynamically adjusts the weights of semantic features according to the spatially consistent relative relationships between objects. It thereby reduces the impact of segmentation errors on the spatial consistency of semantic features.
III-B2 Updating 3D Gaussians
During the mapping process, we assume that the camera pose for the current frame is known. We need to use the current keyframe’s RGBD data to update the GS model of the scene. The updating has two meanings: one is to optimize the established scene parameters, and the other is to generate the new explored GS distribution of the scene.
Following the processes used in Splatam [11] and GS-SLAM [12], silhouette images are rendered to determine the contribution of each surface element (GS) to the map. At the same time, the difference between the projected depth value and the ground truth value of pixels corresponding to newly added GS is checked when they are projected back onto the image plane.
$M\left(p\right)=\left[{Sil\left(p\right)<{T_{s}}}\right]+\left[{\left({{D_{gt}}\left(p\right)-D\left(p\right)}\right)<{T_{d}}}\right]$ (5)
By setting ${\mathbf{d}_{i}}$ in Eq . 3 to a unit value, the silhouette of pixel $p$ is computed. The densification mask $M\left(p\right)$ is calculated according to Eq . 5, where $D\left(p\right)$ represents the depth value of pixel $p$. By calculating the difference between the rendered output from a specific camera position and the ground truth value using Gaussian splatting, the parameters of the 3D Gaussian distribution can be optimized. This problem can also be described as fitting an explicit radiance field to images where the camera pose is known.
After the modeling process discussed in Section III-A, the modeled scene contains three feature channels: spatial position, surface color, and potential semantics. The spatial position and surface color are directly obtained from the RGBD data stream. Meanwhile, the fusion of semantic encoding is supervised by the mask output from a pretrained segmentation model.
${L_{c}}=\lambda{L_{1}}\left({{I_{r}},{I_{gt}}}\right)+\left({1-\lambda}\right)\left[{1-ssim\left({{I_{r}},{I_{gt}}}\right)}\right]$ (6)
The color loss $L_{c}$ is represented as a weighted combination of SSIM [10] and $L1$ loss as in Eq. 6.
${L_{d}}=\sum\limits_{pix}{\left|{D_{pix}^{render}-D_{pix}^{gt}}\right|}$ (7)
The depth loss $L_{d}$ is calculated as in Eq. 7. During the mapping stage, the multi-channel loss is as shown in Eq. 8, where $S_{render}$ represents the rendered semantic feature and $S_{head}$ represents the segmentation mask of the current frame computed by the pretrained model.
${L_{mapping}}={\lambda_{c}}{L_{c}}+{\lambda_{d}}{L_{d}}+{\lambda_{s}}{L_{1}}\left({{S_{render}},{S_{head}}}\right)$ (8)
In Eq. 8, $\lambda_{d}$, $\lambda_{s}$, and $\lambda_{c}$ are predefined hyperparameters used to assign weighted values to the depth, semantic, and color channels respectively.
III-B3 Vitrual Camera Pruning 3D Gaussians
The key aspects of GS-based SLAM are: 1) Distinguishing established high-quality areas from areas that need further optimization, and 2) Identifying and removing outlier points. The former resolves where to add Gaussians, and also plays a key role in camera tracking. Areas of low quality can severely affect the accuracy of pose tracking. The second key aspect resolves where to delete Gaussians. Outlier points will cause holes and defects during image rendering, and these flaws can also affect the accuracy of camera tracking.
The distinction between well-optimized and areas with low quality is implemented through Eq. 5. This section discusses issues related to Gaussians pruning.
MonoGS [13] employs multi-view consistency to eliminate outliers. A Gaussian is deemed visible from a particular view if it is used in rasterization and the ray’s cumulative alpha value has not yet surpassed 0.5. MonoGS maintains a keyframe window, and if Gaussians inserted within the last three keyframes are not observed by at least three other frames in this window, they are considered outliers. Consequently, these Gaussians are removed during the optimization process.
Figure 3: The concept of virtual view pruning for identifying outlier Gaussian points. We analyze only the GS points visible in the current ground-truth view (points $A$, $B$, $C$ in the figure). Point A is not visible from either of the two virtual viewpoints, thus identified as an outlier GS point, and its opacity is degraded during subsequent optimization. While the figure depicts two virtual viewpoints in a planar scenario, our approach creates four virtual cameras by rotating the camera pose from the focal point of each GT view frame along four directions: up, down, left, and right.
Figure 4: Rendered virtual camera views on the ScanNet dataset. he middle images provide a zoomed-in illustration of the effectiveness of Virtual Camera Pruning, where ’vv’ denotes virtual camera view. Eliminating outlier Gaussians not only improves rendering quality but also reduces the storage footprint of the map representation.
Inspired by MonoGS, the proposed NEDS-SLAM introduces a novel Virtual Multi-View Consistency Check approach, as depicted in Fig.3. The points $A$ and $B$ represent outlier GS points, while the GT view denotes the camera pose estimated within the RGBD stream. In the current keyframe, both $A$ and $B$ are visible. However, in the left virtual viewpoint, neither of these outlier points is visible, and in the right virtual viewpoint, only $B$ is visible while $A$ is not. The virtual camera operates alongside the real camera. If a GS point is invisible in all virtual views but visible in the real view, it is then considered an outlier.
The virtual multi-view consistency check method takes advantage of the fast rendering capabilities of the Gaussian Splatting model, enabling the marking of GS points that significantly deviate from the object surface. In subsequent optimization processes, the involvement of outlier GS in the scene is diminished by degrading their opacity. Consistent with [10], GS with near-zero opacity or excessive radius are removed in the mapping process. FSGS [22] used a similar method to improve the quality of images synthesized from novel views. Specifically, FSGS uses the new viewpoint to guide the optimization of grown Gaussians towards a reasonable geometry. Meanwhile, the proposed method in this paper uses multiple synthesized images to eliminate inaccurate Gaussians. As illustrated in Figure. 4, we render virtual views and further optimize the GS distribution only for keyframes. The specific approach for generating virtual views is not fixed. Although Gaussian splatting enables extremely fast virtual view synthesis (nearly 300 FPS), introducing too many viewpoints can compromise the system’s real-time performance. We choose four virtual views along the up, down, left, and right directions, which achieves a desirable balance between effectiveness and efficiency.
III-B4 Camera tracking
The camera tracking phase involves estimating the relative pose of the camera for each new frame, based on the already established map model. The camera pose for the new frame is initialized under the assumption of constant velocity, which includes both a constant linear and angular velocity, as Eq. 9. $T$ and $R$ respectively denote translational and rotational displacement.
$\left[{{T_{t+1}},{R_{t+1}}}\right]=\left[{{T_{t}},{R_{t}}}\right]+\left[{{T_{t}},{R_{t}}}\right]-\left[{{T_{t-1}},{R_{t-1}}}\right]$ (9)
The camera pose is subsequently refined iteratively by minimizing the tracking loss between the ground truth of the color, depth, and semantic channels and the gaussian rendered results from the camera’s perspective.
${L_{tracking}}=\left[{{\lambda_{c}}{L_{c}}+{\lambda_{d}}{L_{d}}+{\lambda_{s}}{L_{1}}\left({{S_{render}},{S_{head}}}\right)}\right]\cdot M$ (10)
$M$ in Eq. 10 is computed as Eq. 5. Artifacts and flaws such as holes and spurious effects caused by outlier gaussians significantly impact the precision of camera tracking. Subsequent experiments demonstrate that the incorporation of semantic loss improve the tracking accuracy. This improvement is attributed to the enriched understanding of the geometric information of objects, facilitated by the integration of semantic features.
Figure 5: The first column shows the RGB reconstruction results. The second column shows the ground truth semantic labels. The third column shows the semantic labels predicted directly on the current frame using M2F [18]. The fourth column shows the semantic reconstruction results using the SGS-SLAM [4] method based on Splatam [11]. The fifth column shows the reconstruction results of our proposed model.
Methods Depth L1[cm] $\downarrow$ LPIPS $\downarrow$ SSIM $\uparrow$ PSNR $\uparrow$ ATE RMSE[cm] $\downarrow$
NICE-SLAM [3] 1.903 0.23 0.81 24.22 2.503
Vox-Fusion [23] 2.913 0.24 0.80 24.41 1.473
Co-SLAM [24] 1.513 0.336 0.94 30.24 1.059
ESLAM [25] 0.945 0.34 0.929 29.08 0.678
Splatam [11] 0.49 0.10 0.97 34.11 0.36
NEDS-SLAM(Ours) 0.47 0.088 0.962 34.76 0.354
Table I: Quantitative comparison of map reconstruction and localization accuracy between our proposed NEDS-SLAM and other NeRF-based dense SLAM methods, averaged over 8 scenes from the Replica dataset.
Methods scene0000 scene0169 scene0181 scene0207 Avg.
NICE-SLAM 12.00 10.90 13.40 6.20 10.63
Vox-Fusion 68.84 27.28 23.30 9.41 32.21
Point-SLAM 10.24 22.16 14.77 9.54 14.18
SplaTAM 12.56 11.09 11.07 7.46 10.54
NEDS-SLAM(Ours) 12.34 11.21 10.35 6.56 10.12
Table II: We quantitatively compare our proposed NEDS-SLAM method against other radiance field-based SLAM on the ScanNet dataset, reporting the tracking metric of camera pose RMSE measured in centimeters.
IV Experiment
IV-A Experimental Setup
Dataset. We evaluate our method on both synthetic and realworld datasets with semantic maps. Following other nerf-based and gaussian-based SLAM methods, for the reconstruction quality, we evaluate quantitatively on 8 synthetic scenes from Replica [26] and qualitatively on 6 scenes from ScanNet [27].
Metrics. We employ several metrics to evaluate the reconstruction quality in our study. These include PSNR, Depth-L1 (on 2D depth maps), SSIM, and LPIPS. Additionally, we assess the accuracy of camera pose estimation using the average absolute trajectory error (ATE RMSE). To evaluate the performance of semantic segmentation, we calculate the mIoU (mean Intersection over Union) score.
Baselines. We compare the tracking and mapping with state-of-the-art methods iMAP, NICE-SLAM, Co-SLAM, ESLAM, and SplaTAM. For semantic segmentation accuracy, we compare with NIDS-SLAM, DNS-SLAM, and SNI-SLAM.
Implementation Details. We conducted experiments using a single NVIDIA 4090 GPU, validating on the REPLICA dataset with the mapping iteration set to 40, tracking iteration set to 60, and SCFF iteration set to 50. After obtaining 384 feature channels through the DINO model, we derived 64-dimensional fused features by applying 2D convolutions separately to the Spatial Features. Finally, we obtained three-dimensional features by passing them through an encoder and embedding them into the GS parameters. We use a learning rate of 0.005 and 0.001 respectively for all learnable parameters on Replica and ScanNet datasets. For camera poses, we only employ a learning rate of 0.0005 in tracking.
Metrics Depth L1[cm] $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ ATE RMSE[cm] $\downarrow$ tracking/frame[s]$\downarrow$ mapping/frame[s] $\downarrow$
Room0 0.31 35.23 0.979 0.082 0.368 0.86 1.52
Room1 0.45 34.86 0.862 0.075 0.403 0.59 1.47
Room2 0.42 35.16 0.983 0.071 0.326 0.67 1.68
Office0 0.31 37.53 0.981 0.091 0.354 0.53 1.71
Office1 0.27 39.71 0.979 0.087 0.284 0.65 1.27
Office2 0.48 32.68 0.973 0.079 0.302 0.60 1.63
Office3 0.58 31.07 0.968 0.103 0.318 0.59 1.47
Office4 0.61 31.82 0.973 0.113 0.473 0.72 1.52
AVG 0.47 34.76 0.962 0.088 0.354 0.65 1.53
Table III: Detailed test results of NEDS-SLAM across 8 scenes from the Replica dataset.
[…]
base model with M2F head 33.57 0.372 26.52
[…]
with SL 32.52 0.368 30.28(+14%)
with FF 33.57 0.371 41.31(+50%)
[…]
Table IV: An ablation study of the NEDS-SLAM, where ’VV’ represents the virtual view pruning method, ’FF’ represents the spatial-consistent semantic fusion method, and ’SL’ represents the utilizing of semantic loss during training.
IV-B Experiment result
Quantitative measures of reconstruction quality using the Replica dataset are presented in Table I. Our method demonstrates competitive performance when compared to other approaches.
The NEDS-SLAM, built upon the foundation of 3DGS, achieves accurate camera localization and semantic reconstruction simultaneously. Table. IV provides a comparison between our method and other neural Implicit approaches in terms of semantic reconstruction performance. Due to the precise representation of object edges offered by the Gaussian radiance field, methods based on neural explicit approaches bring about significant improvements in semantic reconstruction.
Semantic SLAM methods in real-world scenarios often employ a pretrained model to perform semantic segmentation on the RGB-D frames. This semantic information is then utilized in conjunction with geometric information to reconstruct the semantic scene. In an ideal scenario, a perfect semantic segmentation model would exhibit excellent spatio consistency. This means that the same object would be assigned accurate and consistent semantic features across different viewpoints and time instances.
However, the reality often falls short of perfection, and the limitations of semantic SLAM become apparent in the following aspects:
Methods AVG.mIoU[%] $\uparrow$ Room0 Room1 Office0
NIDS-SLAM 82.37 82.45 84.08 85.94
DNS-SLAM 84.77 88.32 84.90 84.66
SNI-SLAM 87.41 88.42 87.43 87.63
Ours 90.78 90.73 91.20 90.42
Table V: A quantitative comparison of our method with existing semantic NeRF-based SLAM approaches on the Replica dataset,For consistency and fair comparison with other methods, we utilize the ground-truth semantic labels from the replica dataset instead of the predicted semantic labels from pre-trained models
  • Inaccurate estimation across frames: Semantic segmentation models may struggle to accurately estimate semantic information across consecutive frames, leading to inconsistencies in the semantic understanding of the scene.
  • Trade-off between performance and inference time: Higher-performing models often require more computational resources and inference time, which can hinder the real-time nature of SLAM systems. The size and complexity of the semantic segmentation model must be carefully balanced to ensure efficient and timely processing.
When testing the dinov2_vit14 model on the replica room0 scene, as shown in Figure. 5, there are noticeable inconsistencies in the predictions for the floor and chairs. This affects the semantic reconstruction quality, but does not impact the reconstruction of the RGB channels. As shown in Figure. 5, NEDS-SLAM effectively filters out the negative impact of spatial semantic inconsistencies, generating robust semantic estimates and providing more accurate semantic reconstruction.
Figure 6: The validation results on the Scannet scene0000_00 dataset. The first row indicates the RGB reconstruction results of NEDS-SLAM, the second row indicates the semantic features predicted by M2F, the third row is the semantic reconstruction results without the Spatially Consistent Feature Fusion (SCFF) module, and the fourth row is the results with the SCFF module.
IV-C Ablation Study
Ablation experiments were designed for the Feature Fusion module and Pruning GS module to test the effectiveness of NEDS-SLAM for semantic reconstruction. As shown in Table. IV, following SGS-SLAM’s approach, we directly incorporated semantic parameters into the GS by calling a pre-trained M2F segmentation model on each RGB frame. The segmented pixel class labels were color-coded and embedded into the GS parameters, resulting in the reconstruction effect displayed in the fourth column of Figure. 5, corresponding to the first row of Table. IV. For the Replica Room0 dataset, the M2F model achieved a semantic segmentation mIoU of 52.4. Employing this method for semantic SLAM resulted in an average mIoU of 26.52, which was used as the baseline approach. After incorporating the virtual view pruning method, outlier GS points that affected reconstruction quality were effectively removed. This led to a 5% increase in PSNR, slightly improved camera pose tracking accuracy, while maintaining the quality of semantic reconstruction. Building upon the baseline, the incorporation of a semantic loss along with depth and RGB losses, weighted at 0.5, 1.0, and 1.0 respectively, resulted in a 14% improvement in semantic reconstruction performance. The addition of the spatial-consistent semantic fusion method, which combined the semantic features from the DinoV2 ViT-14 model with the appearance features from the DepthAnything model, led to nearly a 50% improvement in semantic reconstruction performance. A lightweight encoder-decoder compressed the high-dimensional fused features into a low-dimensional representation, which was then embedded into the GS parameters.
As can be seen in Figure. 6, the semantic features calculated by the M2F model were inconsistent (such as the partitions and books on the table). After processing with the SCFF module, the inconsistencies were resolved and NEDS-SLAM output a more complete semantic reconstruction.
The baseline method suffered from significant semantic noise, resulting in a low semantic reconstruction score. In contrast, the proposed spatial-consistent semantic fusion method effectively addressed the issue of semantic inconsistency, leading to a remarkable improvement in reconstruction performance.
V Conclusion and Limitations
The proposed NEDS-SLAM is an end-to-end semantic SLAM system based on 3D Gaussian splatting. By integrating a Spatially Consistent feature fusion model, NEDS-SLAM effectively addresses the challenges of robustly estimating semantic labels with pre-trained models, significantly enhancing semantic reconstruction performance The proposed Virtual View Pruning method leverages the capabilities of differentiable Gaussian splatting for fast and realistic novel view synthesis. It effectively eliminates outlier GS points in the SLAM process, significantly enhancing the reconstruction quality of the displayed neural radiance fields.
The experiment using public datasets further validated the effectiveness of NEDS-SLAM. However, we also identified existing shortcomings during the experiment. The Virtual view pruning method, which generates virtual views in the keyframe reconstruction pipeline, increases computational load and can affect real-time performance. In the future, we plan to further optimize the virtual view method and take into account semantic reconstruction in dynamic scenes.